lim (3x^2+4/ 5x^4+ 7x^2+1)

x→∞


1
Expert's answer
2021-05-31T07:49:25-0400

We have given the limit


"limx\rightarrow \infty \dfrac{3x^2+4}{5x^4+ 7x^2+1}"


"limx\rightarrow \infty \dfrac{x^2(3+\dfrac{4}{x^2})}{x^4(5+\dfrac{7}{x^2}+\dfrac{1}{x^4})}"


"limx\rightarrow \infty \dfrac{(3+\dfrac{4}{x^2})}{(5+\dfrac{7}{x^2}+\dfrac{1}{x^4})}"


After putting the limit we get 0.


Hence, "limx\rightarrow \infty \dfrac{3x^2+4}{5x^4+ 7x^2+1} = 0"